OpenStudy (anonymous):

Use the binomial expression (p+q)^n to calculate a binomial distribution with n=5 and p=0.3.

OpenStudy (campbell_st):

well the formula I use is basically \[^nC_{r} \times p^x \times (1 - p)^{n - r}\] so you know n and p but not r

OpenStudy (anonymous):

How do i set this up? This is for extra credit, and i really don't understand.

OpenStudy (dan815):

binomial expansion

OpenStudy (anonymous):

Yes

OpenStudy (dan815):

@campbell_st why cant she use q=1-p and cancel everything and say its just 1^5=1 because the odds of P+Q which are complimentary cases happening should that be one anyway

OpenStudy (campbell_st):

you can use q... its just for some people finding q or understanding what it is... can be difficult... so I use (1 - p) they have the same value...

OpenStudy (anonymous):

This is like latin to me you guys.

OpenStudy (campbell_st):

well based on the info provided its \[^5C_{r} (0.3)^r(1 - 0.3)^{5 - r}\]

OpenStudy (dan815):

do what campbell says you can trust him

OpenStudy (anonymous):

The equation in my book is super long and doesn't look like that…

OpenStudy (dan815):

yes queen you have to include all the terms

OpenStudy (anonymous):

I literally wrote the entire question. Thats all the info they gave me.

OpenStudy (dan815):

5C0 p^5 + 5C1 p^4 q +5C2 p^3 q^2 +5C3 p^2 q^3 +5C4 p^1 q^4 +5C5 q^5

OpenStudy (campbell_st):

so for the expansion yon need to let r = 0, 1, 2, 3, 4, 5 so there are 6 terms \[(0.3)^5 + 5 \times (0.3)^4\times(0.7)+ 10 \times (0.3)^3\times (0.7)^2 + 10 \times (0.3)^2 \times (0.7)^3 + 5 \times (0.3) \times (0.7)^4 + (0.7)^5\]

OpenStudy (anonymous):

Thats what it looks like!!! okay!

OpenStudy (dan815):

thats should equal to 1, if im not mistaken let me go check wolfram

OpenStudy (campbell_st):

the missing bit is \[+ 5 \times (0.3) \times(0.7)^4 + (0.7)^5\]

OpenStudy (dan815):

its 1 which is right!

OpenStudy (anonymous):

So my answer is just 1?

OpenStudy (dan815):

i think they were just trying to make you to think that since its (P+q)^n = (p+(1-p)^n = 1^n so you can see it must be 1 no matter what

OpenStudy (anonymous):

Oh.. okay. Well thanks so much!

OpenStudy (dan815):

ok sure, yw!

OpenStudy (anonymous):

Do you have the steps you took to get this answer? I have to type a paragraph (for math i know how dumb) on this and i need to explain that. Thanks Doll.

OpenStudy (anonymous):

@dan815

OpenStudy (dan815):

its a long explaination if you havent seen pascals triangle of this n Choose k stuff

OpenStudy (dan815):

or*

OpenStudy (dan815):

oh wait scratch that okay

OpenStudy (dan815):

the whole point is this binomial distribution, we get to know how much of each case is taking place

OpenStudy (anonymous):

Okay..

OpenStudy (dan815):

like campbell wrote that formula... with 5c1, 5c2, 5c3... each one of those is telling you how this thing is distributed, like the prob of winning twice and failing 3 times or winning 3 times and failing twice and so on

OpenStudy (dan815):

if u plot the terms separating the cases you will see this kind of a graph where in your case you will see 5 lines since n = 5

OpenStudy (dan815):